# Bayesian Optimization of Robustness Measures Using Randomized GP-UCB-based Algorithms under Input Uncertainty

### v1

@@ 5.2 Carrier Lifetime Data
In this section, we conducted experiments using the carrier lifetime value data (Kutsukake et al.,, 2015), which is a measure of the performance of silicon ingots used in solar cells. The original data includes 6586 two-dimensional coordinates on the surface of a silicon ingot and the carrier lifetime ${\rm LT}(x_{1},x_{2})$ at each two-dimensional coordinate $(x_{1},x_{2})$. In this experiment, we focused on $\tilde{\mathcal{X}}\equiv\{(2a+6,2b+6)\mid 1\leq a\leq 88,1\leq b\leq 72\}$, which includes 6336 of these points. We defined the set of design variables $\mathcal{X}$ as a subset of $\tilde{\mathcal{X}}$, and $\mathcal{X}=\{(22a-4,18b-2)\mid 1\leq a\leq 8,1\leq b\leq 8\}$ was used. Also, we used $\Omega=\{(2a-12,2b-10)\mid 1\leq a\leq 11,1\leq b\leq 9\}$ Here, $|\mathcal{X}|=64$, $|\Omega|=99$, $|\mathcal{X}\times\Omega|=6336$, and $\mathcal{X}+\Omega\equiv\{{\bm{x}}+{\bm{w}}\mid{\bm{x}}\in\mathcal{X},{\bm{w}}\in\Omega\}$ is equal to $\tilde{\mathcal{X}}$. For input $(x_{1},x_{2},w_{1},w_{2})\in\mathcal{X}\times\Omega$, the black-box function was defined as $f(x_{1},x_{2},w_{1},w_{2})={\rm LT}(x_{1}+w_{1},x_{2}+w_{2})$. The kernel function of the surrogate model was used as the following Matérn $3/2$ kernel:
@@
where ${\bm{\theta}}=(x_{1}+w_{1},x_{2}+w_{2})$ and ${\bm{\theta}}^{\prime}=(x^{\prime}_{1}+w^{\prime}_{1},x^{\prime}_{2}+w^{\prime}_{2})$. The experiment was performed assuming that there was no observation noise. On the other hand, to stabilize the calculation of the inverse matrix of the kernel matrix, a formal noise variance of $\sigma^{2}_{{\rm noise}}=10^{-6}$ was used when calculating the inverse matrix. The same three robustness measures and nine methods were used as in Section 5.1. The probability mass function used was $p({\bm{w}})=1/99$, and we set $(h,\alpha)=(2.9,4)$. Under this setting, one initial value was chosen randomly, and each algorithm was run until the number of iterations reached 500. This was repeated 100 times, and the average $r_{t}$ for each iteration was calculated. From Figure 3, it can be confirmed that Proposed has the same performance as comparison methods even in the carrier lifetime data.
@@
Figure 3: Average regrets in carrier lifetime data for 100 simulations using each method. The left, middle, and right columns show the results for EXP, PTR, and EXP-MAE, respectively. The length of the error bars represents two times the standard error.
@@ 6 Conclusion
In this study, we proposed the new method for BO of robustness measures for black-box functions with input uncertainty. The proposed method estimates the optimal solution using posterior means, samples the parameters of GP-UCB from the probability distribution, and determines the next evaluation point based on the estimated solution, credible intervals, and posterior variance. In Section 4, we provided upper bounds for the expected value of the regret and cumulative regret, and showed that their orders for commonly used robustness measures, including the expectation measure, are $O(\sqrt{\gamma_{t}/t})$ and $O(\sqrt{t\gamma_{t}})$, respectively.
Compared to existing methods, the proposed method has the following three advantages. First, the proposed method is not specialized for a specific robustness measure, as in the methods in Beland and Nair, (2017); Iwazaki et al., 2021a , but can be used in a general way for any robustness measure that satisfies the condition in (3.2). Next, compared to the method in Inatsu et al., 2024a , which is not specialized for a particular robustness measure like the proposed method, it has the advantage of not requiring hyper-parameters for the acquisition function. Finally, in this study, we derived the order of the expected value of the regret and cumulative regret defined using the estimated solution based on the posterior mean. To the best of our knowledge, this is the first study to provide these regret bounds for various robustness measures.
On the other hand, disadvantages of the proposed method are as follows. First, the most significant feature of the proposed method is that it randomly replaces the parameters of GP-UCB, but this alone does not significantly improve practical performance. Next, as the derivation of the theoretical analysis of the proposed method, we derived the high-probability bound, but compared to the usual GP-UCB-based bound, the fact that $\delta^{-1}$ appears in the bound and is not tight is also a problem. Finally, it is generally difficult to calculate the best index, $\hat{t}$, among the estimated optimal solutions $\hat{\bm{x}}_{t}$.
Improving the above disadvantages is one of the important directions for future study. In addition, extensions that include cases where the input space is high-dimensional or there are multiple robustness measures are left as future work.
@@ A.1 Extension to Continuous Settings when $\mathcal{X}$ is Continuous and $\Omega$ is Finite
Let $\mathcal{X}$ be a continuous set, and let $\Omega$ be a finite set. Suppose that $\mathcal{X}$ is a compact and convex set with $\mathcal{X}\subset[0,r]^{d_{1}}$. In this setting, the only difference between Algorithm 1 and an extension of the proposed method is the distribution of $\beta_{t}$. Specifically, by using $\kappa^{(1)}_{t}$ with $1\leq\kappa^{(1)}_{1}\leq\kappa^{(1)}_{2}\leq\cdots$, we define $\beta_{t}=2\log\kappa^{(1)}_{t}+\xi_{t}$. Theoretically valid values of $\kappa^{(1)}_{1},\ldots,\kappa^{(1)}_{t}$ and the theoretical analysis are given in Appendix A.1.1. The pseudo-code for the case when $\mathcal{X}$ is continuous and $\Omega$ is finite is given in Algorithm 2.
Algorithm 2 RRGP-UCB for robustness measures when $\mathcal{X}$ is continuous and $\Omega$ is finite
@@
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (3.4)
   Select the next evaluation point $\bm{w}_{t}$ by (3.5)
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
@@
In order to derive the theoretical guarantee, we introduce additional two assumptions.
@@
There exists a non-decreasing and concave function $q_{1}(a)$ such that $q_{1}(0)=0$ and
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f({\bm{x}}^{\prime},{\bm{w}}))|\leq q_{1}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f({\bm{x}}^{\prime},{\bm{w}})|\right)\quad{\rm for}\ {\bm{x}},{\bm{x}}^{\prime}\in\mathcal{X}.$
For Assumption A.1, note that when ${\bm{w}}$ is fixed, $f({\bm{x}},{\bm{w}})$ is a GP on $\mathcal{X}$. In GP-based BOs, similar assumptions to Assumption A.1 are used in, for example, Srinivas et al., (2010); Takeno et al., (2023). Here, if the target robustness measure is described in Table 4, for Assumption A.2, we can use $q(a)$ in Table 4 as $q_{1}(a)$. Details are described in Appendix C.5. Then, the following theorem holds.
@@
Assume that Regularity assumption, Assumption 4.1, A.1 and A.2, and (3.2) hold. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(1)}_{t}=(1+\lceil b_{1}d_{1}rt^{2}(\sqrt{\log(a_{1}d_{1}|\Omega|)}+\sqrt{\pi}/2)\rceil^{d_{1}})|\Omega|$, and define $\beta_{t}=2\log\kappa^{(1)}_{t}+\xi_{t}$. Let $q_{1}(a)$ and $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.2 and 4.1, respectively. Then, the following holds if Algorithm 2 is performed:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
@@
Under the assumption in Theorem A.1, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
@@
Under the assumption in Theorem A.1, the following holds:
@@
where $\hat{t}$ is given by (4.3), and functions $q_{1}(\cdot)$, $h_{i}(\cdot)$ and each coefficient are given in Theorem A.1. Moreover, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@ A.2 Extension to Continuous Settings when $\mathcal{X}$ is Finite and $\Omega$ is Continuous
$G^{\dagger}_{t-1}({\bm{x}})=\{g({\bm{x}},{\bm{w}})\mid^{\forall}{\bm{w}}\in\Omega,g({\bm{x}},{\bm{w}})\in Q^{\dagger}_{t-1}({\bm{x}},{\bm{w}})\},$
@@
${\rm lcb}^{\dagger}_{t-1}({\bm{x}})\leq\inf_{g({\bm{x}},{\bm{w}})\in G^{\dagger}_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}})),\ \sup_{g({\bm{x}},{\bm{w}})\in G^{\dagger}_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}}))\leq{\rm ucb}^{\dagger}_{t-1}({\bm{x}}).$ (A.1)
For commonly used robustness measures described in Table 3, we can use ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ in Table 3 as ${\rm lcb}^{\dagger}_{t-1}({\bm{x}})$ and ${\rm ucb}^{\dagger}_{t-1}({\bm{x}})$, respectively. Using this, we define the credible interval $Q^{\dagger}_{t-1}({\bm{x}})=[{\rm lcb}^{\dagger}_{t-1}({\bm{x}}),{\rm ucb}^{\dagger}_{t-1}({\bm{x}})]$. Here, we emphasize that $Q^{\dagger}_{t-1}({\bm{x}})$ is the credible interval for $\rho(f({\bm{x}},[{\bm{w}}]_{t}))\equiv F^{\dagger}_{t}({\bm{x}})$, not $F({\bm{x}})$. We define the estimated solution $\hat{\bm{x}}^{\dagger}_{t}$ by using $\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$ as follows:
$\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}})).$ (A.2)
@@
where $\tilde{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm ucb}^{\dagger}_{t}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}^{\dagger}_{t}({\bm{x}}))_{+}$.
@@
The pseudo-code for the proposed method is given in Algorithm 3.
Algorithm 3 RRGP-UCB for robustness measures when $\mathcal{X}$ is finite and $\Omega$ is continuous
@@
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}^{\dagger}_{t}$ by $\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (A.3)
   Select the next evaluation point $\bm{w}_{t}$ by ${\bm{w}}_{t}=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{w}})$
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
@@
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f({\bm{x}},[{\bm{w}}]_{t}))|\leq q_{2}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f({\bm{x}},[{\bm{w}}]_{t})|\right)$
@@
For Assumption A.3, if ${\bm{x}}$ is fixed, then $f({\bm{x}},{\bm{w}})$ is a GP on $\Omega$. Hence, as in the case of Assumption A.1, we can obtain a sufficient condition for Assumption A.3 by using the derivative of the kernel function. Furthermore, for Assumption A.4, by letting $f({\bm{x}},[{\bm{w}}]_{t})\equiv f^{\dagger}({\bm{x}},{\bm{w}})$, we can use $q(a)$ in Table 4 as $q_{2}(a)$ if the target robustness measure is described in Table 4. Details are given in Appendix C.5. In addition, for the optimal solution ${\bm{x}}^{\ast}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}F({\bm{x}})$ and estimated solution $\hat{\bm{x}}^{\dagger}_{t}$, we define the regret $r^{\dagger}_{t}$ and cumulative regret $R^{\dagger}_{t}$ as follows:
@@
As in the case of $q(a)$, commonly used robustness measures described in Table 4, we can use $q(a)$ in Table 4 as $q^{\dagger}(a)$. Then, the following theorem holds.

### v2

@@ 5.3 Carrier Lifetime Data
In this section, we conducted experiments using the carrier lifetime dataset (Kutsukake et al.,, 2015), which quantifies the performance of silicon ingots used in solar cells. The original dataset includes 6586 two-dimensional coordinates on the surface of a silicon ingot and the corresponding carrier lifetime values, denoted by ${\rm LT}(x_{1},x_{2})$ at each coordinate $(x_{1},x_{2})$. In this experiment, we focused on the subset $\tilde{\mathcal{X}}\equiv\{(2a+6,2b+6)\mid 1\leq a\leq 88,1\leq b\leq 72\}$, which includes 6336 of these points. The set of design variables $\mathcal{X}$ was defined as a subset of $\tilde{\mathcal{X}}$, specifically $\mathcal{X}=\{(22a-4,18b-2)\mid 1\leq a\leq 8,1\leq b\leq 8\}$. In addition, we defined $\Omega=\{(2a-12,2b-10)\mid 1\leq a\leq 11,1\leq b\leq 9\}$. This results in $|\mathcal{X}|=64$, $|\Omega|=99$, and $|\mathcal{X}\times\Omega|=6336$, with the set $\mathcal{X}+\Omega\equiv\{{\bm{x}}+{\bm{w}}\mid{\bm{x}}\in\mathcal{X},{\bm{w}}\in\Omega\}$ equal to $\tilde{\mathcal{X}}$. For each input $(x_{1},x_{2},w_{1},w_{2})\in\mathcal{X}\times\Omega$, the black-box function was defined as $f(x_{1},x_{2},w_{1},w_{2})={\rm LT}(x_{1}+w_{1},x_{2}+w_{2})$. The kernel function used in the surrogate model was the Matérn $3/2$ kernel, defined as follows:
@@
where ${\bm{\theta}}=(x_{1}+w_{1},x_{2}+w_{2})$ and ${\bm{\theta}}^{\prime}=(x^{\prime}_{1}+w^{\prime}_{1},x^{\prime}_{2}+w^{\prime}_{2})$. The experiment was performed under the assumption of no observation noise. However, to ensure numerical stability when computing the inverse of the kernel matrix, a nominal noise variance of $\sigma^{2}_{{\rm noise}}=10^{-6}$ was added. The same three robustness measures and nine methods as described in Section 5.1 were employed. The probability mass function was set to $p({\bm{w}})=1/99$, and the parameters $(h,\alpha)=(2.9,4)$. Under this setting, one initial point was selected at random, and each algorithm was run for 500 iterations. This procedure was repeated 100 times, and the average regret $r_{t}$ was computed for each iteration. As shown in Figure 4, the Proposed method demonstrated performance comparable to that of the baseline methods, even on the carrier lifetime dataset.
@@
Figure 4: Average regrets on carrier lifetime dataset over 100 simulations for each method. Left, middle, and right columns correspond to results for EXP, PTR, and EXP-MAE, respectively. Since BBBMOBO (fixed) and Proposed (fixed) are the same in EXP, only Proposed (fixed) is displayed in the figures on the left column. Error bars represent twice standard error.
@@ 6 Conclusion
In this paper, we proposed a new method for BO of robustness measures for black-box functions with input uncertainty. The proposed method estimated the optimal solution using posterior means, sampled the parameters of GP-UCB from a probability distribution, and determined the next evaluation point based on the estimated solution, credible intervals, and posterior variance. In Section 4, we provided upper bounds on the expected regret and cumulative regret and showed that their orders for commonly used robustness measures, including the expectation measure, were $O(\sqrt{\gamma_{t}/t})$ and $O(\sqrt{t\gamma_{t}})$, respectively.
Compared to existing methods, the proposed method offered the following three advantages. First, unlike the methods in Beland and Nair, (2017); Iwazaki et al., 2021a , which were tailored to specific robustness measures, our method was applicable more generally to any robustness measure satisfying the condition in (3.2). Second, in contrast to the method in Inatsu et al., 2024a , which was also not restricted to a particular robustness measure, our method did not require hyperparameters for the acquisition function. Third, we derived the order of the expected regret and cumulative regret defined in terms of the estimated solution based on the posterior mean. To the best of our knowledge, this study is the first to establish expected (cumulative) regret bounds for various robustness measures.
However, the proposed method also had certain limitations. Most significantly, while it randomly replaced the parameters of GP-UCB, a key feature of the method, this mechanism alone did not significantly improve practical performance. Additionally, although we derived a high-probability bound for the theoretical analysis, the appearance of $\delta^{-1}$ in the bound, which was not tight compared to typical GP-UCB-based bounds with respect to $\delta$, posed a limitation. Furthermore, although the proposed method can be extended to continuous spaces, it has the disadvantage of losing the advantages of the proposed method, namely, that $\beta_{t}$ does not need to be adjusted and that no increase in $t$ is required. Finally, it was generally difficult to calculate the best index $\hat{t}$ among the estimated optimal solutions $\hat{\bm{x}}_{t}$.
Addressing these disadvantages remains an important direction for future study.
@@ A.1 Extension to Continuous Settings when $\mathcal{X}$ is Continuous and $\Omega$ is Finite
Let $\mathcal{X}$ be a continuous set, and let $\Omega$ be a finite set. Suppose that $\mathcal{X}$ is a compact and convex set with $\mathcal{X}\subset[0,r]^{d_{1}}$. In this setting, the only difference between Algorithm 1 and an extension of the proposed method is the distribution of $\beta_{t}$. Specifically, by using $\kappa^{(1)}_{t}$ with $1\leq\kappa^{(1)}_{1}\leq\kappa^{(1)}_{2}\leq\cdots$, we define $\beta_{t}=2\log\kappa^{(1)}_{t}+\xi_{t}$. Theoretically valid values of $\kappa^{(1)}_{1},\ldots,\kappa^{(1)}_{t}$ and the theoretical analysis are given in Appendix A.1.1. The pseudo-code for the case when $\mathcal{X}$ is continuous and $\Omega$ is finite is provided in Algorithm 3.
Algorithm 3 RRGP-UCB for robustness measures when $\mathcal{X}$ is continuous and $\Omega$ is finite.
@@
   Generate $\xi_{t}$ from chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho({\mu_{t-1}({\bm{x}},\cdot)})$
   Select next evaluation point $\bm{x}_{t}$ by (3.4)
   Select next evaluation point $\bm{w}_{t}$ by (3.5)
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding observed data
@@
To derive the theoretical guarantee, we introduce additional two assumptions.
@@
There exists a non-decreasing, concave function $q_{1}(a)$ such that $q_{1}(0)=0$ and
$|\rho({f({\bm{x}},\cdot)})-\rho({f({\bm{x}}^{\prime},\cdot)})|\leq q_{1}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f({\bm{x}}^{\prime},{\bm{w}})|\right)\quad{\rm for}\ {\bm{x}},{\bm{x}}^{\prime}\in\mathcal{X}.$
For Assumption A.1, note that when ${\bm{w}}$ is fixed, $f({\bm{x}},{\bm{w}})$ is a GP on $\mathcal{X}$. In GP-based BOs, similar assumptions to Assumption A.1 are used in, for example, Srinivas et al., (2010); Takeno et al., (2023). Here, for the Bayes risk (expectation), worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures described in Table 4 in Inatsu et al., 2024a , we can use $q^{(m)}$ in the table as $q_{1}(a)$ in Assumption A.2. Details are described in Appendix C.5. Then, the following theorem holds.
@@
Assume that the regularity assumption, Assumptions 4.1, A.1 and A.2, and (3.2) hold. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(1)}_{t}=(1+\lceil b_{1}d_{1}rt^{2}(\sqrt{\log(a_{1}d_{1}|\Omega|)}+\sqrt{\pi}/2)\rceil^{d_{1}})|\Omega|$, and define $\beta_{t}=2\log\kappa^{(1)}_{t}+\xi_{t}$. Let $q_{1}(a)$ and $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumptions A.2 and 4.1, respectively. Then, if Algorithm 3 is performed, the following holds:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken over all sources of randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
@@
Under the assumptions of Theorem A.1, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for the other measures.
@@
Under the assumptions of Theorem A.1, the following holds:
@@
where $\hat{t}$ is given by (4.4), and functions $q_{1}(\cdot)$, $h_{i}(\cdot)$ and all coefficients are as defined in Theorem A.1. Moreover, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@ A.2 Extension to Continuous Settings when $\mathcal{X}$ is Finite and $\Omega$ is Continuous
$G^{\dagger}_{t-1}({\bm{x}})=\{{g({\bm{x}},\cdot)}\mid{\rm for\ all}\ {\bm{w}}\in\Omega,g({\bm{x}},{\bm{w}})\in Q^{\dagger}_{t-1}({\bm{x}},{\bm{w}})\},$
@@
${\rm lcb}^{\dagger}_{t-1}({\bm{x}})\leq\inf_{{g({\bm{x}},\cdot)}\in G^{\dagger}_{t-1}({\bm{x}})}\rho({g({\bm{x}},\cdot)}),\ \sup_{{g({\bm{x}},\cdot)}\in G^{\dagger}_{t-1}({\bm{x}})}\rho({g({\bm{x}},\cdot)})\leq{\rm ucb}^{\dagger}_{t-1}({\bm{x}}).$ (A.1)
For commonly used robustness measures described in Table 3 in Inatsu et al., 2024a , we can use ${\rm lcb}^{(m)}_{t-1}({\bm{x}})$ and ${\rm ucb}^{(m)}_{t-1}({\bm{x}})$ in the table as ${\rm lcb}^{\dagger}_{t-1}({\bm{x}})$ and ${\rm ucb}^{\dagger}_{t-1}({\bm{x}})$, respectively. Using this, we define the credible interval $Q^{\dagger}_{t-1}({\bm{x}})=[{\rm lcb}^{\dagger}_{t-1}({\bm{x}}),{\rm ucb}^{\dagger}_{t-1}({\bm{x}})]$. Here, we emphasize that $Q^{\dagger}_{t-1}({\bm{x}})$ is the credible interval for $\rho({f^{\dagger}_{t}({\bm{x}},\cdot)})\equiv F^{\dagger}_{t}({\bm{x}})$, not $F({\bm{x}})$, where $f^{\dagger}_{t}({\bm{x}},{\bm{w}})\equiv f({\bm{x}},[{\bm{w}}]_{t})$. We define the estimated solution $\hat{\bm{x}}^{\dagger}_{t}$ by using ${\mu^{\dagger}_{t-1}({\bm{x}},\cdot)}$ as follows:
$\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho({\mu^{\dagger}_{t-1}({\bm{x}},\cdot)}).$ (A.2)
@@
where $\tilde{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm ucb}^{\dagger}_{{t-1}}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}^{\dagger}_{{t-1}}({\bm{x}}))_{+}$.
@@
The pseudo-code for the proposed method is provided in Algorithm 4.
Algorithm 4 RRGP-UCB for robustness measures when $\mathcal{X}$ is finite and $\Omega$ is continuous.
@@
   Generate $\xi_{t}$ from chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}^{\dagger}_{t}$ by $\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho({\mu^{\dagger}_{t-1}({\bm{x}},\cdot)})$
   Select next evaluation point $\bm{x}_{t}$ by (A.3)
   Select next evaluation point $\bm{w}_{t}$ by ${\bm{w}}_{t}=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{w}})$
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding observed data
@@
$|\rho({f({\bm{x}},\cdot)})-\rho({f^{\dagger}_{t}({\bm{x}},\cdot)})|\leq q_{2}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f({\bm{x}},[{\bm{w}}]_{t})|\right)$
@@
For Assumption A.3, if ${\bm{x}}$ is fixed, then $f({\bm{x}},{\bm{w}})$ is a GP on $\Omega$. Hence, as in the case of Assumption A.1, we can obtain a sufficient condition for Assumption A.3 by using the derivative of the kernel function. Furthermore, for Assumption A.4, we can use $q^{(m)}(a)$ in Table 4 in Inatsu et al., 2024a as $q_{2}(a)$ if the target robustness measure is the Bayes risk (expectation), worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, or mean absolute deviation measure described in the table. Details are given in Appendix C.5. In addition, for the optimal solution ${\bm{x}}^{\ast}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}F({\bm{x}})$ and estimated solution $\hat{\bm{x}}^{\dagger}_{t}$, we define the regret $r^{\dagger}_{t}$ and cumulative regret $R^{\dagger}_{t}$ as follows:
@@
As in the case of $q(a)$, for the Bayes risk (expectation), worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures described in Table 4 in Inatsu et al., 2024a , we can use $q^{(m)}(a)$ in the table as $q^{\dagger}(a)$. Then, the following theorem holds.